Espresso Express operates a number of espresso coffee stands in busy suburban malls. The fixed weekly expense of a coffee stand is $600 and the variable cost per cup of coffee served is $0.75. Required: 1. Fill in the following table with your estimates of the company's total cost and average cost per cup of coffee at the indicated levels of activity. 2. Does the average cost per cup of coffee served increase, decrease, or remain the same as the number of cups of coffee served in a week increases?

Complete this question by entering your answers in the tabs below.

Fill in the following table with your estimates of the company's total cost and average cost per cup of coffee at the indicated levels of activity. (Round off the average cost per cup of coffee to the nearest tenth of a cent.)

Cups of Coffee Served in a Week
1,200 1,300 1,400
Fixed cost
Variable cost
Total cost $0 $0 $0
Average cost per cup of coffee served
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Espresso Express Note
Cups of Coffee Served in a Week 1200 1300 1400 A
Fixed cost       600.00       600.00       600.00 B
Variable cost per unit           0.75           0.75           0.75 C
Variable cost       900.00       975.00 1,050.00 D=C*A
Total cost 1,500.00 1,575.00 1,650.00 E=B+D
Average cost per cup of coffee served           2.50           2.63           2.75 G=E/A